the difference between the circumference and the radius of a circle is 37cm .find its area.
1
Expert's answer
2012-11-27T09:01:31-0500
L-R=2piR-R=37
So, R=37/(2pi-1)
AreaS=piR^2=pi(37/(2pi-1))^2=154.085